The length of a rectangle is 5 inches more than it's width the perimeter is 66 inches what is the dimension of the rectangle
Answer by TimothyLamb(4379)

L = w + 5
2L + 2w = 66
---
2L + 2w = 66
2(w + 5) + 2w = 66
2w + 10 + 2w = 66
4w = 56
---
w = 14
L = 19
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php